OpenStudy (kitkat1):

Medal for best answer!!! Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 4−x and y = 2x + 3 intersect are the solutions of the equation 4−x = 2x + 3. Part B: Make tables to find the solution to 4−x = 2x + 3. Take the integer values of x between −3 and 3. Part C: How can you solve the equation 4−x = 2x + 3 graphically?

OpenStudy (kitkat1):

@foolofatook

OpenStudy (foolofatook):

part a makes no sense.

OpenStudy (foolofatook):

im not good with graphs

OpenStudy (kitkat1):

Can u answer the 2nd part?

OpenStudy (kitkat1):

I was mostly confused with the 2nd part

OpenStudy (foolofatook):

ok

OpenStudy (foolofatook):

lets start with plugging in the lowest number and go up

OpenStudy (kitkat1):

4-(1)=2(1)+3 3=5 false 4-(2)=2(2)+3 2=7 false 4-(3)=2(3)+3 1=9 false 4-(4)=2(4)+3 4=3 false 4-(5)=2(5)+3 5=1 false 4-(6)=2(6)+3 6=1 false 4-(7)=2(7)+3 7=-3 false

OpenStudy (kitkat1):

this is what i did but dont know how that will help me find the answer

OpenStudy (foolofatook):

the question says the values are between -3 and 3

OpenStudy (foolofatook):

4−-3 = 2*-3 + 3.

OpenStudy (kitkat1):

4-(-3)=2(-3)+3 7=3 false

OpenStudy (anonymous):

A. We have two lines: y = 2-x and y = 4x+3 Given two simultaneous equations that are both required to be true.. the solution is the points where the lines cross... Which is where the two equations are equal.. Thus the solution that works for both equations is when 2-x = 4x+3 because where that is true is where the two lines will cross and that is the common point that satisfies both equations.

OpenStudy (foolofatook):

srry doing two things at once

OpenStudy (foolofatook):

4−-2 = 2*-2 + 3 ?

OpenStudy (kitkat1):

@kaitlynfisher You made a mistake it is actually y=4-x and y=2x+3

OpenStudy (anonymous):

Solving it graphically: You just insert the functions and see where they cross:

OpenStudy (kitkat1):

4-(2)=2(2)+3 2=7 false

OpenStudy (foolofatook):

4−-1 = 2*-1 + 3.

OpenStudy (kitkat1):

5=1 false

OpenStudy (foolofatook):

4-0=2*0+3.?

OpenStudy (kitkat1):

4=3 false

OpenStudy (foolofatook):

4−1 = 2*1 + 3.

OpenStudy (kitkat1):

3=5 false

OpenStudy (foolofatook):

4−2 = 2*2 + 3.

OpenStudy (kitkat1):

2=7 false

OpenStudy (kitkat1):

How is this going to help me find the answer?

OpenStudy (foolofatook):

it seems there are no whole number answers

OpenStudy (kitkat1):

Then what is the answer?

OpenStudy (anonymous):

The answer is \[x=\frac{1}{3}\] So it doesnt work trying with whole numbers.

OpenStudy (kitkat1):

how do I explain that?

OpenStudy (kitkat1):

Okay it is x=1/3 how can that be explained?

OpenStudy (foolofatook):

write it as a decimal, 0.33 repeating, wich means it has a line over the threes

OpenStudy (foolofatook):

4−(0.33) = 2(0.33) + 3.

OpenStudy (anonymous):

You see from this table, that is would be between 0 and 1. Then you would start taking like 1/4, 2/4, 3/4 and then you would find that it is between 1/4 and 2/4, this way you could find the answer using a table.

OpenStudy (kitkat1):

3.67=3.66 false

OpenStudy (kitkat1):

Okay whatever what about the part C?

OpenStudy (foolofatook):

this stupid graph... i dont know @Tommynaiter ?

OpenStudy (kitkat1):

Okay well I have 2 more questions I will give u a medal now and post it in a different thread can u come and help me with them?

OpenStudy (foolofatook):

sure. tag me

OpenStudy (kitkat1):

Okay